(1 point) The times for the mile run of a large group of male college students are approximately Normal with mean 7.02 minutes and standard deviation 0.72 minutes. Use the 68-95-99.7 rule to answer the following questions. (Start by making a sketch of the density curve you can use to mark areas on.) (a) what range of times covers the middle 99.7% of this distribution? (b) What percent of these men run a mile in less than 8.46 minutes? (a) from minutes to minutes(1 point) The heights of women aged 20 to 29 follow approximately the N(64, 2.67) distribution. Men the same age have heights distributed as N(69.3, 2.87). What percent of young women are taller than the mean height of young men?(1 point) Changing the mean and standard deviation of a Normal distribution by a moderate amount can greatly change the percent of observation in the tails. Suppose that a college is looking for applicants with SAT math scores 720 and above (a) In 2007, the scores of men on the math SAT followed the N(533, 116) distribution. What percent of men scored 720 or better? (b) Womens SAT math scores that year had the N(499, 110) distribution. What percent of women scored 720 or better?

Please help with these questions. Will give thumbs up top rating! thank you
Answer #1

a)accordang to central limit theorem the middle of 99.7 % covers 3 standard deviations.

Hence from 7.02-0.72*3 to 7.02+0.72*3

= 4.86 to 9.18

b) at 8.46 the Standard deviation count is 8.46-7.02=1.44 /0.72=2

Since at 2 standard deviation in an normal distribution it covers 97.5 %.

Next part

Since the distribution of women height follows a normal distribution with mean =64 and standard deviation as 2.67 , hence as asked the what percentage of women follows more height than men mean height.

Mean height of men is 69.3 so bg using. Z statistics

Z= \frac{X-\mu}{\sigma}

Z at X= 69.3 is 1.985 calculated hence or 2 hence here according to central limit theorem 2.357% of women height are more than mean men height.

Again,

a) If SAT distribution follow a normal distribution with mean =533 and standard deviation as 116 hence by using Z statistic formula

Z= \frac{X-\mu}{\sigma}

Z=1.612 now calculating the are more than at Z =1.612 by Z table or by calculator as 0.0534 means 5.34% age of men having scores 720 or more than 720.

b. Again if women SAT scores follows normal distribution hence again by Z statistic formula

Z= \frac{X-\mu}{\sigma}

Z at 720 calculated as 2.009 or 2 hence by central limit theorem 2.227 % of women have 720 or better.